# Dynamics on invariant tori emerging through forced symmetry breaking in phase oscillator networks

### v1

@@ 5.1. Dynamics on the perturbation of $\mathrm{SDD}$
The small interval of multistability due to the existence of bifurcating secondary branches leads to new coexisting stable synchrony patterns. Figure 3 shows the phases of the oscillators along one period of these secondary stable periodic orbits. Shown are the phases of the six oscillators against the rescaled integration time for $\delta\approx 0.403722$, that is within the interval between a saddle-node and a torus bifurcation where the periodic orbits on the isolas on figure 2 are all stable. The color indicates the deviation from $\pi$ of $\theta_{\sigma,k}$, where black means $\theta_{\sigma,k}=\pi$ and white means $\theta_{\sigma,k}=0$ or $2\pi$. Here, Panels (a) and (b) correspond to stable periodic orbits lying on top and bottom of the secondary branch that emanates from $\psi=\pi$, respectively and Panels (c) and (d) show the same for the the secondary branch emanating from $\psi=0$. This shows that in these new synchrony patterns the phases on ‘$\mathrm{D}$’ populations are not exactly $\pi$ apart, but rather wiggle around $\pi$ and even get close for extremely short periods of time.
@@
Figure 5. Evolution of periodic orbits along the overlapping principal saddle-node curves $\mathsf{SN}_{p}$ in figure 4. Panels (a)–(d) correspond to the points $r_{a},r_{b},r_{c},r_{d}$, respectively. Each panel shows (1) a time series of the coordinate $\theta_{2,1}$ with respect to the rescaled Auto integration time for orbits on each principal saddle-node curve in green and red, (2) a relative position between the two ‘$\mathrm{D}$’ populations on each principal saddle-node curve in green and red, and (3) how each ‘$\mathrm{D}$’ populations deviate from $\pi$ as $(r,\delta)$ moves away from the origin.
The saddle-node curve $\mathsf{SN}_{p}$ in figure 4(a) corresponds to a saddle node bifurcation of the primary branches of periodic orbits $\gamma_{\frac{\pi}{2}}$ (dark blue) and $\gamma_{\pi}$ (dark red) as $r$ is decreased at a value of $r$ that is $\mathcal{O}(\delta^{3/2})$; cf. Figure 4(b). (The same occurs with the branches $\gamma_{\frac{3\pi}{2}}$ (dark blue) and $\gamma_{0}$ (dark red) at numerically the same value.) Starting with $r=0.2$ and decreasing $r$ while keeping $\delta=0.01$ fixed, the saddle-node bifurcation occurs at $r=r_{a}\approx 0.0011$. At these bifurcation points, the phases of the ‘$\mathrm{D}$’ populations of are $\frac{3\pi}{4}$ and $\frac{7\pi}{4}$ apart, which is half way between the relative positions of the `$\mathrm{D}$’ populations initially. This saddle-node bifurcation is a consequence of the $\mathbb{Z}_{2}$ symmetry inherited from reduced dynamics, since by decreasing $r$ towards zero we are decreasing the second harmonic in the symmetry breaking perturbation that allow the solutions to exist in the first place. This saddle-node bifurcation on the primary branches can be then continued in $(r,\delta)$ tracing out the curve $\mathsf{SN}_{p}$. The curve converges to $(0,0)$ in parameter plane; solutions along $\mathsf{SN}_{p}$ at $r_{a}\approx 0.0011$, $r_{b}=0.01$ (panel (b)), $r_{c}=0.02$, and $r_{d}=0.03$ are shown in Figure 5. Naturally, the curve $\mathsf{SN}_{p}$ bounds the existence of other bifurcations $\mathsf{SN}_{1}$$\mathsf{SN}_{5}$ along the secondary branches since the primary branches $\gamma_{\psi^{*}}$ with $\psi^{*}\in\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\}$ emerge there.
@@ 5.2. Dynamics on the perturbation of $\mathrm{SSD}$
with $\varphi=\phi_{2}-\phi_{1}$. The variable $\phi_{3}$ has constant motion and the nullclines of $\varphi$ provide periodic orbits in system (48). The quantity $\rho=\rho(r,\alpha,\beta)=\frac{\cos\alpha}{2r\cos(2\beta)}$ determines the influence of second harmonics in the $\varphi$-dynamics unless $r=0$ or $\beta=\beta^{*}$ with $\cos(2\beta^{*})=0$. For $\lvert\rho\rvert>1$ there is one stable and one unstable periodic orbit located at $\phi=0,\pi$; which of these orbits is stable/unstable depends on the sign of $\rho$. The periodic orbit $\varphi=\pi$ undergoes a pitchfork bifurcation at $\lvert\rho\rvert=1$ and two symmetry-related periodic orbits $\varphi=\varphi^{*}$ and $\varphi=2\pi-\varphi^{*}$ emanate from there and persist for $\lvert\rho\rvert<1$, where $\varphi^{*}$ is solution of $\cos\alpha+2r\cos\varphi\cos(2\beta)=0$. The critical value $\lvert\rho\rvert=1$ is a threshold for the onset of the influence of second harmonics in the dynamics of (48). For $\lvert\rho\rvert<1$ system (48) has two stable and two unstable periodic orbits and the phase space looks similar to figure 1, which is the situation for the parameter values given by (51). Note that the choice of $\phi$ that allows to obtain the reduced system (52) has a different interpretation than the variable $\psi$ for $\mathrm{SDD}$: Rather than the relative position between desynchronized populations, $\varphi$ corresponds the the relative position between the synchronized (‘$\mathrm{S}$’) populations on $\mathrm{SSD}$.
Figure 6. Bifurcation diagram of periodic orbits of the full system (19) with the perturbation driven by (22) and (25) for $\delta>0$ for the perturbed torus $\mathrm{SSD}$. The rest of the parameters are as in figure 2. Panel (a) shows four branches of hyperbolic periodic orbits $\gamma_{0},\gamma_{\frac{\pi}{2}},\gamma_{\pi},\gamma_{\frac{3\pi}{2}}$ emanating from $\psi^{*}\in\big{\{}0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\big{\}}$ at $\delta=0$. The vertical green line located at $\delta=0$ represents the unpertubed case, in which the torus $\mathrm{SSD}$ is foliated by neutrally stable periodic orbits. Panel (b) highlights the isola of secondary branches of periodic orbits emanating from the branch $\gamma_{\pi}$, and Panel (c) shows a zoom of the origin of the branch of secondary periodic orbits that emanates from $\gamma_{0}$.
@@ 6. Discussion
We analyzed the dynamics on invariant tori that arise by perturbing relative equilibria in the context of a network of coupled oscillators. First, we showed that the parameterization method developed in [15] can be applied to relative equilibria of a continuous $\mathbb{T}^{m}$ symmetry. This yields explicit equations for the perturbed tori as well as the dynamics thereon to any order. Second, we computed the first-order approximation of the torus and its dynamics explicitly by solving the corresponding conjugacy equations. Third, we used the first-order approximation to generate starting data for numerical continuation in Auto. This allowed to compute branches of periodic orbits that emerge on the perturbed tori as the coupling strength is increased beyond the regime where the first-order approximation is valid (or even beyond the existence of an invariant torus). The choice of explicitly parameterized phase interaction functions allows to link the results to parameters in physical models, e.g., through phase reduction [22, 26, 27].
@@
which exists for an open set of parameters $(q,K)$. The work done in this paper can be used in the larger problem of studying global features of system (15) that relate to perturbations of the heteroclinic structure. We can now use the knowledge gathered about the local properties of the tori $\mathrm{SSD}$, $\mathrm{SDD}$, etc. and use it to study how the connections between these sets are affected under forced symmetry breaking: One would expect that trajectories close to the stable manifold of a perturbed saddle torus $\mathbf{T}$ first approach the torus, then move towards a saddle periodic orbit $\gamma$ on $\mathbf{T}$ (at a timescale determined by the perturbation parameter $\delta$), before leaving the neighborhood of $\mathbf{T}$ near the unstable manifold of $\gamma$. Thus, the unstable manifold of $\gamma$ (within the unstable manifold of $\mathbf{T}$) gives information about the global dynamics of the perturbed system. While this is a question for further research, the numerical setup implemented here—using continuation of orbit segments in a boundary-value-problem setup—can be also used to tackle some of these global questions.
Ackowledgements
@@
Here, $u=(u_{1},\ldots,u_{6})\in\mathbb{T}^{6}$, $F=F_{\delta}:\mathbb{T}^{6}\times\mathbb{R}^{k}\to\mathbb{\mathbb{T}}^{6}$ is the right-hand side of (19) and $\lambda\in\mathbb{R}^{k}$ is its vector of parameters, including the perturbation parameter $\delta$. Importantly, $u:[0,1]\to\mathbb{T}^{6}$ so that any orbit segment is parameterized over the unit interval $[0,1]$; the actual integration time $T$ is considered as a separate parameter. The function $u$ is a unique solution of (53) if suitable boundary conditions are imposed at one or both end points $u(0)$ and $u(1)$. Therefore, each orbit segment is defined in terms of the conditions one imposes upon $u(0)$ and $u(1)$.
The usual boundary condition for a periodic in $\mathbb{R}^{n}$ is $u(1)-u(0)=0$. This is already implemented in the Auto routines for the continuation of a periodic orbit in parameters when we set the corresponding problem type by choosing IPS=2 in the Auto-constants file; see [14]. In the context of coupled oscillators, since the phase variable $\theta$ (and therefore its rescaled version $u$) lies on a torus a periodic orbit can be defined by
@@
Here $\eta=(\eta_{1},\eta_{2},\eta_{3})$ is also a vector of internal parameters that is used to monitor when a solution to (55) with initial condition $v(0)=v^{0}$ is periodic in $\mathbb{T}^{3}$, for suitable fixed $(l_{1},l_{2},l_{3})\in\mathbb{Z}^{3}$. We start the continuation of (55)–(57) setting $v^{1}=v^{0}$ and with the integration time $T$$v^{1}$ and $\eta$ free during the continuation, so that the problem is well-posed. After some initial exploration, we can detect values for $l_{1},l_{2}$ and $l_{3}$; for instance, on $\mathrm{SDD}$ we have $l_{1}=2$, $l_{2}=l_{3}=-2$. We stop the continuation when $\eta=0$, thus obtaining an approximation of a periodic orbit in $\mathbb{T}^{3}$. Once we have this initial periodic orbit, we can embed it in $\mathbb{T}^{6}$ as described in section 5 in order to get an initial periodic orbit that is a solution to (53), (54) in $\mathbb{T}^{6}$. The values of $k_{1}$$k_{6}$ in (54) are obtained through the embedding as well.
@@
Since the system on $\mathbb{T}^{6}$ still has a continuous $\mathbb{T}$ symmetry that leads to lack of hyperbolicity, we actually consider the reduced system on $\mathbb{T}^{5}$. Specifically, we set $\theta_{\sigma,k}\to\theta_{\sigma,k}-\theta_{1,1}$, that is, we consider phase differences of each of the oscillators with respect to $\theta_{1,1}$. This is the actual system numerically studied in section 5 with the setup explained above, with the corresponding changes considering its dimension.

### v2

@@ 5.1. Dynamics on the perturbation of $\mathrm{SDD}$
The small interval of multistability due to the existence of bifurcating secondary branches leads to new coexisting stable synchrony patterns. Figure 3 shows the phases of the oscillators along one period of these secondary stable periodic orbits. Shown are the phases of the six oscillators against the rescaled integration time for $\delta\approx 0.403722$, that is within the interval between a saddle-node and a torus bifurcation where the periodic orbits on the isolas on Figure 2 are all stable. The color indicates the deviation from $\pi$ of $\theta_{\sigma,k}$, where black means $\theta_{\sigma,k}=\pi$ and white means $\theta_{\sigma,k}=0$ or $2\pi$. Here, Panels (a) and (b) correspond to stable periodic orbits lying on top and bottom of the secondary branch that emanates from $\psi=\pi$, respectively and Panels (c) and (d) show the same for the the secondary branch emanating from $\psi=0$. This shows that in these new synchrony patterns the phases on $\mathrm{D}$’ populations are not exactly $\pi$ apart, but rather wiggle around $\pi$ and even get close for extremely short periods of time.
@@
Figure 5. Evolution of periodic orbits along the overlapping principal saddle-node curves $\mathsf{SN}_{p}$ in Figure 4. Panels (a)–(d) correspond to the points $r_{a},r_{b},r_{c},r_{d}$, respectively. Each panel shows (1) a time series of the coordinate $\theta_{2,1}$ with respect to the rescaled Auto integration time for orbits on each principal saddle-node curve in green and red, (2) a relative position between the two $\mathrm{D}$’ populations on each principal saddle-node curve in green and red, and (3) how each $\mathrm{D}$’ populations deviate from $\pi$ as $(r,\delta)$ moves away from the origin.
The saddle-node curve $\mathsf{SN}_{p}$ in Figure 4(a) corresponds to a saddle node bifurcation of the primary branches of periodic orbits $\gamma_{\frac{\pi}{2}}$ (dark blue) and $\gamma_{\pi}$ (dark red) as $r$ is decreased at a value of $r$ that is $\mathcal{O}(\delta^{3/2})$; cf. Figure 4(b). (The same occurs with the branches $\gamma_{\frac{3\pi}{2}}$ (dark blue) and $\gamma_{0}$ (dark red) at numerically the same value.) Starting with $r=0.2$ and decreasing $r$ while keeping $\delta=0.01$ fixed, the saddle-node bifurcation occurs at $r=r_{a}\approx 0.0011$. At these bifurcation points, the phases of the $\mathrm{D}$’ populations of are $\frac{3\pi}{4}$ and $\frac{7\pi}{4}$ apart, which is half way between the relative positions of the $\mathrm{D}$’ populations initially. This saddle-node bifurcation is a consequence of the $\mathbb{Z}_{2}$ symmetry inherited from reduced dynamics, since by decreasing $r$ towards zero we are decreasing the second harmonic in the symmetry breaking perturbation that allow the solutions to exist in the first place. This saddle-node bifurcation on the primary branches can be then continued in $(r,\delta)$ tracing out the curve $\mathsf{SN}_{p}$. The curve converges to $(0,0)$ in parameter plane; solutions along $\mathsf{SN}_{p}$ at $r_{a}\approx 0.0011$, $r_{b}=0.01$ (panel (b)), $r_{c}=0.02$, and $r_{d}=0.03$ are shown in Figure 5. Naturally, the curve $\mathsf{SN}_{p}$ bounds the existence of other bifurcations $\mathsf{SN}_{1}$$\mathsf{SN}_{5}$ along the secondary branches since the primary branches $\gamma_{\psi^{*}}$ with $\psi^{*}\in\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\}$ emerge there.
@@ 5.2. Dynamics on the perturbation of $\mathrm{SSD}$
with $\varphi=\phi_{2}-\phi_{1}$. The variable $\phi_{3}$ has constant motion and the nullclines of $\varphi$ provide periodic orbits in system (49). The quantity $\rho=\rho(r,\alpha,\beta)=\frac{\cos\alpha}{2r\cos(2\beta)}$ determines the influence of second harmonics in the $\varphi$-dynamics unless $r=0$ or $\beta=\beta^{*}$ with $\cos(2\beta^{*})=0$. For $\lvert\rho\rvert>1$ there is one stable and one unstable periodic orbit located at $\phi=0,\pi$; which of these orbits is stable/unstable depends on the sign of $\rho$. The periodic orbit $\varphi=\pi$ undergoes a pitchfork bifurcation at $\lvert\rho\rvert=1$ and two symmetry-related periodic orbits $\varphi=\varphi^{*}$ and $\varphi=2\pi-\varphi^{*}$ emanate from there and persist for $\lvert\rho\rvert<1$, where $\varphi^{*}$ is solution of $\cos\alpha+2r\cos\varphi\cos(2\beta)=0$. The critical value $\lvert\rho\rvert=1$ is a threshold for the onset of the influence of second harmonics in the dynamics of (49). For $\lvert\rho\rvert<1$ system (49) has two stable and two unstable periodic orbits and the phase space looks similar to Figure 1, which is the situation for the parameter values given by (52). Note that the choice of $\varphi$ in the reduced system (53) has a different interpretation than the variable $\psi$ for $\mathrm{SDD}$: Rather than the relative position between desynchronized populations, $\varphi$ corresponds the the relative position between the synchronized ($\mathrm{S}$’) populations on $\mathrm{SSD}$.
Figure 6. Bifurcations of periodic orbits of the full system (16) with the perturbation driven by (19) and (26) for $\delta>0$ for the perturbed torus $\mathrm{SSD}$. The rest of the parameters are as in Figure 2. Panel (a) shows four branches of hyperbolic periodic orbits $\gamma_{0},\gamma_{\frac{\pi}{2}},\gamma_{\pi},\gamma_{\frac{3\pi}{2}}$ emanating from $\psi^{*}\in\big\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\big\}$ at $\delta=0$. The vertical green line located at $\delta=0$ represents the unperturbed case, in which the torus $\mathrm{SSD}$ is foliated by neutrally stable periodic orbits. Panel (b) highlights the isola of secondary branches of periodic orbits emanating from the branch $\gamma_{\pi}$, and Panel (c) shows a zoom of the origin of the branch of secondary periodic orbits that emanates from $\gamma_{0}$.
@@ 6. Discussion
We analyzed the dynamics on invariant tori that arise by perturbing relative equilibria in the context of a network of coupled oscillators. First, we showed that the parameterization method developed in [15] can be applied to relative equilibria of a continuous $\mathbb{T}^{m}$ symmetry. This yields explicit equations for the perturbed tori as well as the dynamics thereon to any order. Second, we computed the first-order approximation of the torus and its dynamics explicitly by solving the corresponding conjugacy equations. Third, we used the first-order approximation to generate starting data for numerical continuation in Auto. This allowed to compute branches of periodic orbits that emerge on the perturbed tori as the coupling strength is increased beyond the regime where the first-order approximation is valid (or even beyond the existence of an invariant torus). The choice of explicitly parameterized phase interaction functions yields a link to parameters in physical models, e.g., through phase reduction [23, 26, 27].
@@
which exists for an open set of parameters $(q,K)$. The work done in this paper can be used in the larger problem of studying global features of system (16) that relate to perturbations of the heteroclinic structure. We can now use the knowledge gathered about the local properties of the tori $\mathrm{SSD}$, $\mathrm{SDD}$, etc. and use it to study how the connections between these sets are affected under forced symmetry breaking: One would expect that trajectories close to the stable manifold of a perturbed saddle torus $\mathbf{T}$ first approach the torus, then move towards a saddle periodic orbit $\gamma$ on $\mathbf{T}$ (at a timescale determined by the perturbation parameter $\delta$), before leaving the neighborhood of $\mathbf{T}$ near the unstable manifold of $\gamma$. Thus, the unstable manifold of $\gamma$ (within the unstable manifold of $\mathbf{T}$) gives information about the global dynamics of the perturbed system. While this is a question for further research, the numerical setup implemented here—using continuation of orbit segments in a boundary-value-problem setup—can be also used, together with state-of-the-art techniques to compute invariant manifolds of saddle invariant objects, to tackle some of these global questions. For an entry point on these continuation methods we refer the reader to [28].
@@
Here, $u=(u_{1},\dotsc,u_{6})\in\mathbb{T}^{6}$, $F=F_{\delta}:\mathbb{T}^{6}\times\mathbb{R}^{k}\to\mathbb{\mathbb{T}}^{6}$ is the right-hand side of (16) and $\lambda\in\mathbb{R}^{k}$ is its vector of parameters, including the perturbation parameter $\delta$. Importantly, $u:[0,1]\to\mathbb{T}^{6}$ so that any orbit segment is parameterized over the unit interval $[0,1]$; the actual integration time $T$ is considered as a separate parameter. The function $u$ is a unique solution of (54) if suitable boundary conditions are imposed at one or both end points $u(0)$ and $u(1)$. Therefore, each orbit segment is defined in terms of the conditions one imposes upon $u(0)$ and $u(1)$.
The usual boundary condition for a periodic in $\mathbb{R}^{n}$ is u(1)u(0)=0u(1)-u(0)=0. This is already implemented in the Auto routines for the continuation of a periodic orbit in parameters when we set the corresponding problem type by choosing IPS=2 in the Auto-constants file; see [14]. In the context of coupled oscillators, since the phase variable $\theta$ (and therefore its rescaled version $u$) lies on a torus a periodic orbit can be defined by
@@
[…]
@@
Note that the full system on $\mathbb{T}^{6}$ still has a continuous $\mathbb{T}$ symmetry that acts by a common phase shift to all oscillators. Since this leads to lack of hyperbolicity, we consider in practice the reduced system on $\mathbb{T}^{5}$. Specifically, we set $\theta_{\sigma,k}\to\theta_{\sigma,k}-\theta_{1,1}$, that is, we consider phase differences of each of the oscillators with respect to $\theta_{1,1}$. This is the actual system numerically studied in Section 5 with the setup explained above, with the corresponding changes considering its dimension. Here, in the boundary conditions (55) defining a periodic orbit, the information from $k_{1}$ is passed to the rest of the constants via $k_{i}\to k_{i}-k_{1}$ for $i=2,\dotsc,5$.